A rose garden has 21 bushes in one row, 17 bushes in the next row, then 13, and so on. If there are 66 bushes in the garden, in how many rows are they planted?


Answer by edjones(8007)

Aritmetic sequence:
c=a[1]-d
c=21-(-4)=25
.
a[n]=dn+c
a[n]=-4n+25
.
(n/2)(a[1]+a[n])=S[n]
(n/2)(21-4n+25)=66
n(46-4n)=132
46n-4n^2=132
4n^2-46n+132=0
2n^2-23n+66=0
(n-6)(2n-11)=0
n=6 rows
.
Ed